OpenStudy (itsmichelle29):

Given the parent functions f(x) = 3x + 2 and g(x) = 5x − 10, what is g(x) − f(x)? g(x) − f(x) = 5x − 8 − 3x g(x) − f(x) = 5x − 12 − 3x g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x + 12 − 5x

jimthompson5910 (jim_thompson5910):

\[\Large \color{red}{f(x) = 3x+2}\] \[\Large \color{blue}{g(x) = 5x-10}\] --------------------------------------------- \[\Large \color{blue}{g(x)}-\color{red}{f(x)}=\color{blue}{(5x-10)}-\color{red}{(3x+2)}\] I'll let you finish

OpenStudy (itsmichelle29):

What next

OpenStudy (itsmichelle29):

@jim_thompson5910

OpenStudy (itsmichelle29):

i get =2x−12

OpenStudy (itsmichelle29):

and that not the answer

jimthompson5910 (jim_thompson5910):

look through the choices and see which ones have a "-12" in it for some reason, they didn't combine the x terms, idk why

OpenStudy (itsmichelle29):

g(x) = 5^x − 10,

jimthompson5910 (jim_thompson5910):

oh these are exponentials?

OpenStudy (itsmichelle29):

5 is to the power of x sorry

jimthompson5910 (jim_thompson5910):

I see now

OpenStudy (itsmichelle29):

3 is not only 5

jimthompson5910 (jim_thompson5910):

you cannot do \[\Large 5^x - 3^x = 2^x\]

jimthompson5910 (jim_thompson5910):

\[\Large a^x - b^x \ne (a-b)^x\]

OpenStudy (itsmichelle29):

so whats the answer

jimthompson5910 (jim_thompson5910):

you tell me

OpenStudy (itsmichelle29):

B i think or A

jimthompson5910 (jim_thompson5910):

what's your reasoning on each?

OpenStudy (itsmichelle29):

B for -12 or A for -10+2

jimthompson5910 (jim_thompson5910):

but you have to distribute the negative through

jimthompson5910 (jim_thompson5910):

\[\Large \color{red}{f(x) = 3^x+2}\] \[\Large \color{blue}{g(x) = 5^x-10}\] --------------------------------------------- \[\Large \color{blue}{g(x)}-\color{red}{f(x)}=\color{blue}{(5^x-10)}-\color{red}{(3^x+2)}\] \[\Large g(x) - f(x) = 5^x - 10 - 3^x - 2\] \[\Large g(x) - f(x) = 5^x - 12 - 3^x\]

jimthompson5910 (jim_thompson5910):

notice how \(\Large -( 3^x + 2)\) becomes \(\Large - 3^x - 2\)

OpenStudy (itsmichelle29):

Oh okayyyy can u help me with one more

jimthompson5910 (jim_thompson5910):

ok one more

OpenStudy (itsmichelle29):

Given the parent functions f(x) = log10 x and g(x) = 3x − 1, what is f(x) • g(x)? f(x) • g(x) = log10 (3x − 1)x f(x) • g(x) = log10 x3x − 1 f(x) • g(x) = 3x log10 x + log10 x f(x) • g(x) = log10 x − 3x log10 x

OpenStudy (itsmichelle29):

i think its A

OpenStudy (itsmichelle29):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

can you post a screenshot of this?

OpenStudy (itsmichelle29):

yeah

OpenStudy (itsmichelle29):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

none of the answer choices are correct. There's one that's close, but it's off by a negative sign

jimthompson5910 (jim_thompson5910):

there has to be a typo somewhere

jimthompson5910 (jim_thompson5910):

oh nvm I see the answer

OpenStudy (itsmichelle29):

pkay thanks anyways

jimthompson5910 (jim_thompson5910):

hint: y*log(x) = log(x^y)

OpenStudy (itsmichelle29):

if fine i turned the tyest in

OpenStudy (itsmichelle29):

thanks